determine whether the graphs of the pairs of lines are parallel
x+8=y
y-x= -4

Found 2 solutions by ikleyn, fractalier:
Answer by ikleyn(52946)

.
Yes, they are parallel.

Both have the slope 1.


Answer by fractalier(6550)

If you solve them for y, that is, put them in y = mx + b form, you can verify that their slopes are indeed the same...thus the lines are parallel...
So from
x+8=y
y-x= -4
we get
y = x + 8 and
y = x - 4
See, parallel, both having a slope of one.